OpenStudy (anonymous):

Solve the equation, 2^(x+1)-2^(x-1)+2^(x-3)-2^(x-4)=50.

OpenStudy (dumbcow):

expand \[2*2^{x} - \frac{1}{2}2^{x}+\frac{1}{8}2^{x}-\frac{1}{16}2^{x} = 50\] \[2^{x}(\frac{25}{16}) = 50\] \[2^{x} = 32\] \[x=5\]

OpenStudy (anonymous):

Hey dumbcow may u explain why 2^x(25/16)=50 become 2^x=32.

OpenStudy (dumbcow):

multiply by (16/25) on both sides

OpenStudy (anonymous):

thanks